Solve the equation show steps to get there
X^2 = x + 6

Answer by mananth(16946)

x^2=x+6
x^2-x-6=0
x^2-3x+2x-6=0
x(x-3)+2(x-3)=0
(x-3)(x+2)=0
x=3 OR -2